# Similarity-Navigated Conformal Prediction for Graph Neural Networks

### v1

@@ Appendix B Algorithm overview
7:  $\forall k\in\mathcal{V}$, compute $\hat{s}(\boldsymbol{x}_{n+1},k)$
8:  return  $\mathcal{C}_{\alpha}(\boldsymbol{x}_{n+1})=\{k|\hat{s}(\boldsymbol{x}_{n+1},k)\leq\hat{q}\}$
@@ B.1 Time Complexity
Constructing the similarity graph will have particularly high time complexity for large-scale graphs if we compute the feature similarity between each node pair, i.e., time complexity is $\mathcal{O}(N^{2})$. To address this problem, we randomly select a subset of nodes as the target node set with a size of $M\ll N$. We then calculate the similarity between the nodes in calibration and testing sets and the nodes in the target set, and select $k$-NN. This approach reduces the time complexity to $\mathcal{O}(MN)$.
Appendix C Detailed Results
@@
In this section, we report the detailed experimental results. Firstly, we report the results of $\mathrm{Coverage}$, $\mathrm{Size}$, $\mathrm{SH}$ across different datasets using different GNNs models in Table 4, Table 5, Table 6, Table 7. Then, we report the results of $\mathrm{Coverage}$, $\mathrm{Size}$, $\mathrm{SH}$ using RAPS as the basic non-conformity scores in Table 8. Finally, We added two additional datasets CiteSeer and Amazon Computers to observe changes in the distribution of mean scores for each label in Table LABEL:fig:citeseer-size and Table LABEL:fig:com-size. The results show that SNAPS achieve superior performance.
@@
Appendix D Experiments on Image Classification
To verify that our method could be adapted to image classification problems, we test SNAPS on ImageNet and various models. We split the test dataset containing 50000 images into 25000 images for the calibration set and 25000 images for the test set. The models are pretrained Imagenet classifiers from the torchvision repository [Paszke et al., 2019] with standard normalization, resize, and crop parameters. Moreover, all models are calibrated by the Temperature scaling procedure [Guo et al., 2017a] on the calibration set. All experiments are conducted with ten trials, and the mean results are reported.
@@ Appendix D Experiments on Image Classification
Appendix E More Details on Dataset and Models

### v2

@@ Appendix B Algorithm overview
7:  $\forall k\in\mathcal{V}$, compute $\boldsymbol{\hat{{S}}}_{n+1,k}=\hat{s}(s(\boldsymbol{x}_{n+1},k),\boldsymbol{A}_{s})$
8:  return  $\mathcal{C}_{\alpha}(\boldsymbol{x}_{n+1})=\{k|\boldsymbol{\hat{{S}}}_{n+1,k}\leq\hat{q}\}$
@@ B.1 Time Complexity
The time complexity of SNAPS is primarily determined by the computation of corrected scores. In this work, we use one-hop nodes and nodes with high feature similarity to correct the ego node. In the transductive setting, this complexity applies to the entire graph. Consequently, one-hop generalization requires $\mathcal{O}(E)$ runtime, while $k$-NN generalization requires $\mathcal{O}(NM)$, where $E$ is the number of edges, $N$ is the number of test nodes, $M$ is the number of nodes sampled to correct the scores of test nodes, with $M\ll N$ for large graphs. For example, we randomly sample $M=80,000$ nodes from the original set of 2,449,029 nodes in the OGBN Products dataset. Finally, the time complexity of SNAPS is $\mathcal{O}(E+NM)$.
Time complexity of $k$-NN.
Calculating pairwise similarities is inherently parallelizable, which significantly enhances the efficiency of $k$-NN. Furthermore, there have been some approximation methods that could be used to significantly speed up the computation for large graphs, such as NN-Descent [Dong et al., 2011] that can be easily implemented under MapReduce, achieving an approximate $k$-NN graph in $\mathcal{O}(N^{1.14})$ empirically.
Appendix C Additional Experiments and Experimental Details
C.1 Experiments on heterophilous graph datasets
To analyze the performance of SNAPS on heterophilous networks, we conduct experiments on two common heterophilous graph datasets. Detailed statistics of two datasets are shown in Appendix F.
We choose FSGNN [Maurya et al., 2022] as the GNN model. We adopt the dataset splits of Geom-GCN [Pei et al., 2020], i.e. splitting nodes into $60\%/20\%/20\%$ for training/validation/testing. To evaluate the performance of the CP methods, we divide the test set equally into the calibration and evaluation sets. For DAPS, we set $\lambda=0.5$. For SNAPS, we set $\lambda=0.5,\mu=0$, and $k=20$, i.e., neglecting the structural neighborhood. To construct a cosine similarity-based $k$-NN graph for two heterophilous datasets, we utilize the embeddings from FSGNN. In Table 5, empirical results at $\alpha=0.1$ and $0.15$ are presented. The experimental results show that SNAPS consistently outperforms the baselines on heterophilous networks.
Table 5: Results of $\mathrm{Coverage}$, $\mathrm{Size}$ on two heterophilous graph datasets. We report the average calculated from FSGNN with 100 conformal splits at different significance levels ($\alpha=0.05,0.1$). Bold numbers indicate optimal performance.
$\alpha=0.1$ $\alpha=0.15$
[…]
APS DAPS SNAPS APS DAPS SNAPS APS DAPS SNAPS APS DAPS SNAPS
Chameleon 78.09±0.93 0.904 0.905 0.904 1.95 2.75 1.70 0.850 0.853 0.851 1.62 2.23 1.32
Squirrel 73.72±2.19 0.900 0.897 0.900 2.41 3.05 2.27 0.851 0.848 0.850 1.89 2.37 1.64
Discussion on heterophilous graph datasets.
In this work, we utilize the embeddings from FSGNN to construct the similarity graph. Although this serves as a simple workaround for heterophilous graphs, it does not effectively address the core issue in heterophily, where the expectation is that the embeddings would effectively capture the heterophilous networks. Therefore, we plan to construct the similarity graph based solely on graph structure and node features to enhance the performance of SNAPS in future work.
C.2 Comparison with CF-GNN
To compare with CF-GNN, we randomly select 20 nodes per class for training/validation, setting the calibration set size to 1,000 [Huang et al., 2023b]. The APS scores are used as the basic non-conformity scores. The average results are calculated from 10 GCN with each trial of 100 conformal splits. Moreover, we introduce a new metric, i.e., $\mathrm{Time}$, to evaluate the running time for each trial. As shown in Table 6, SNAPS outperforms CF-GNN in both metrics.
Table 6: Results of $\mathrm{Size}$, $\mathrm{Time}$ on three graph datasets for CF-GNN and SNAPS. We report the results at different significance levels ($\alpha=0.05,0.1$). Bold numbers indicate optimal performance.
Size at $\alpha=0.05$ Size at $\alpha=0.1$ Time
Datasets CF-GNN SNAPS CF-GNN SNPAS CF-GNN SNAPS
CoraML 2.60 1.68 1.68 1.31 142s 0.73s
PubMed 2.13 1.62 1.86 1.35 148s 0.98s
CiteSeer 3.07 1.84 1.96 1.39 124s 0.71s
Appendix D Detailed Results
@@
In this section, we report the detailed experimental results. Firstly, we report the results of $\mathrm{Coverage}$, $\mathrm{Size}$, $\mathrm{SH}$ across different datasets using different GNNs models in Table 7, Table 8, Table 9, Table 10. Then, we report the results of $\mathrm{Coverage}$, $\mathrm{Size}$, $\mathrm{SH}$ using RAPS as the basic non-conformity scores in Table 11. Finally, we added two additional datasets CiteSeer and Amazon Computers to observe changes in the distribution of mean scores for each label in Figure LABEL:fig:citeseer-size and Figure LABEL:fig:com-size. The results show that SNAPS achieves superior performance.
@@
Appendix E Experiments on Image Classification
To verify that our method could be adapted to image classification problems, we test SNAPS on ImageNet and various models. We split the test dataset containing 50000 images into 25000 images for the calibration set and 25000 images for the test set. The models are pretrained Imagenet classifiers from the torchvision repository [Paszke et al., 2019] with standard normalization, resize, and crop parameters. Moreover, all models are calibrated by the Temperature scaling procedure [Guo et al., 2017] on the calibration set. All experiments are conducted with ten trials, and the mean results are reported.
@@ Appendix E Experiments on Image Classification
Appendix F More Details on Dataset and Models